when you have 20=5f+15 what would you do to get the 5f onto the other side of the = sign? Would it be a positive or a negative?
Answer by richwmiller(17219)

don't know why you want to do that
20=5f+15
subtract 15 from both sides
5=5f
1=f
but anyways
20=5f+15
20-5f+15
-5f=-5
f=1
negative
you subtract 5f from both sides
you want 0 instead of 5f so you have to subtract
if you added 5f you get
20+5f=5f+5f+15
20+5f=10f+15